OpenStudy (anonymous):

A 100 degree arc of a circle has a length of 7cm. To the nearest centimeter, what's the radius of the circle? Please help emergency! Medal/fan

OpenStudy (anonymous):

The arc length of a circle is defined as follows: $$ s = \theta r $$ You have s and theta, now just algebraically solve for r

OpenStudy (anonymous):

But what's s in this problem

OpenStudy (anonymous):

s is the arc length :)

OpenStudy (anonymous):

Oh thanks!!! Are u up for one more?

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

Mr. H races on a bike with 26 inch diameter wheels. When he is traveling at a speed of 44 ft/sec, how many revolutions per min are his wheels making?

OpenStudy (anonymous):

Okay, let's first get everything in similar units. It will be easier to work in inches, since it will give us whole numbers. 44 ft = how many inches?

OpenStudy (anonymous):

528 inches

OpenStudy (anonymous):

Bueno. So every second, Mr. H is traveling 528 inches. This is where arc length is useful!

OpenStudy (anonymous):

The arc length is just like saying, if you laid the rubber of the tire out as a line, how long would it be? 

OpenStudy (anonymous):

How does that apply here?

OpenStudy (anonymous):

Well, we need to find the arc length of the tire in order to find out how many rotations per second he is making

OpenStudy (anonymous):

So, we know that $$ s = \theta r $$ We need the radius. Do you know how to get the radius from the diameter?

OpenStudy (anonymous):

ya 26/2 = 13

OpenStudy (anonymous):

Perfect, so the diameter is 13 in. Now, how many degrees are in a full circle?

OpenStudy (anonymous):

360

OpenStudy (anonymous):

Great. So the arc length is?

OpenStudy (anonymous):

whats theta in this case